OpenStudy (anonymous):

sim[plify the squeare root of 24

OpenStudy (reemii):

do you know that - (for \(a>0\)), \(\sqrt{a^2}=a\) - (for \(a>0\)), \(\sqrt{ab}=\sqrt{a}\sqrt b\) ? If so, write the number 24 as a product of two numbers, one of which is a perfect square. and then you're almost done, use the properties I wrote above.

OpenStudy (reemii):

what's your answer?